Problem 3:

The traffic manager at ABC Electronics needs to ship 20,000 lbs. of a popular TV from the main warehouse in Chicago to ABC’s regional warehouse in Denver (1,500 miles). The load will fit in a single 40-ft trailer. The traffic manager is using a contract carrier to perform the shipment. The carrier has quoted to the following pricing information:

• Transportation line-haul cost is US$15.50 per mile in the carrier’s 40-ft trailer

. • The pickup and delivery costs for the shipment is US$250.00.

Based on the data in problem 3, the carrier has informed the traffic planner that if ABC could assemble a second truck with the same weight of 20,000 lbs., the line-haul cost could be cut to US$3.75 and the pickup and delivery cost for both trucks cut to US$175. The traffic planner is interested in the proposal. However, the large quantity would mean that one of the trucks would have to make a terminal stop in Kansas City before proceeding on to Denver. A terminal stop would cost US$450.00. What would the cost of the proposal be to ABC?

What is the total line-haul cost for the shipment?
Answer #1

Answer:

Information in the question:

Distance from Chicago to Denver = 1500 miles

Proposal A – 40 feet trailer:

Line Haul charges/ mile (US$) = US$ 15.5 / mile

Total Line Haul charges from Chicago to Denver = US $ 15.5 x 1500 miles = US$ 23,250

Pick up & delivery charges (US$) = US$ 250

Total cost of proposal of Proposal A = 23250+250 = US$ 23500.00

Proposal B – 2 trucks:

Line Haul charges/ mile / truck (US$) = US$ 3.75

Total Line Haul charges for 2 truck (US$) = 2 x 3.75 x 1500 = US$ 11,250

Pick up & delivery charges (US$) for 2 trucks = US$ 175

Terminal stop charge at Denver for one of the truck (US$) = US$ 450

Total cost of proposal of Proposal B = 11250+175+450 = US$ 11875.00